OpenStudy (anonymous):

Write the given sentence as an equation. Tim’s age in 7 years will be three times what it was 19 years ago. A. 3(t + 7) = t – 19 B. t + 19 = 3(t – 7) C. t + 7 = 3(t – 19) D. 3(t + 19) = t – 7

hartnn (hartnn):

19 years ago = t-19 in 7 years = t+7 so can u get it now ?

OpenStudy (anonymous):

so it either b or c

hartnn (hartnn):

no, not b., b doesn't have t-19

OpenStudy (anonymous):

Not B, it would be (t+19) not minus

OpenStudy (anonymous):

i mean A or c

hartnn (hartnn):

three times what it was 19 years ago.--->> 3(t-19)

OpenStudy (anonymous):

If you're trying to get 3 times what the age was 19 years ago, the 3 would be attached to (t-19)

OpenStudy (anonymous):

Thannks you guys!!! idk who to give best responce

OpenStudy (anonymous):

Give it to hartnn, he had it first :P

hartnn (hartnn):

m not here for medals...